OpenStudy (anonymous):

IM DOING applications with polynomials::: 1:) Solve: x4 − 5x3 + 11x2 − 25x + 30 = 0 2:) Solve the equation: x2 + 5x + 22 = 0 3:) Solve: x3 − 2x2 + 10x + 136 = 0

OpenStudy (anonymous):

please~~~

OpenStudy (anonymous):

It kind like complicated to write down all the work, and i just online by my cell phone. I did your works. Do you want i give you the problem after I factored?

OpenStudy (anonymous):

please ! i need the answer mainly

OpenStudy (anonymous):

Ok 1. (x-2)(x-3)(x^2+5)=0

OpenStudy (anonymous):

Im going to do 2.

OpenStudy (anonymous):

2. Use quadratic formula:x= (-5+,-3isqrt (7))/2

OpenStudy (anonymous):

These are the answer choices for # 2 0 -2.5 15.75 No solution

jimthompson5910 (jim_thompson5910):

the discriminant of #2 is D = b^2 - 4ac D = (5)^2 - 4(1)(22) D = 25 - 88 D = -63 a negative discriminant means what?

OpenStudy (anonymous):

So #2 it just has the imagine solution, it is not real solution for it.------>no solution

OpenStudy (anonymous):

These are the answer chocies for #3 -4, 3 − 5i, 3 + 5i -2, 3 − 5i, 3 + 5i -4, 8 No solution

jimthompson5910 (jim_thompson5910):

x^3 - 2x^2 + 10x + 136 = 0 (-2)^3 - 2(-2)^2 + 10(-2) + 136 = 0 -8 - 2(4) + 10(-2) + 136 = 0 -8 - 8 - 20 + 136 = 0 100 = 0 which is FALSE, so x = -2 is NOT a root. so that rules out choice B

OpenStudy (anonymous):

alright so?

jimthompson5910 (jim_thompson5910):

x^3 - 2x^2 + 10x + 136 = 0 (-4)^3 - 2(-4)^2 + 10(-4) + 136 = 0 -64 - 2(16) + 10(-4) + 136 = 0 -64 - 32 - 40 + 136 = 0 0 = 0 which is TRUE, so x = -4 is a root. that means that the answer is either A or C

jimthompson5910 (jim_thompson5910):

plug in x = 8 and see if it makes the polynomial zero

OpenStudy (anonymous):

it does

jimthompson5910 (jim_thompson5910):

no it doesn't, try it again

OpenStudy (anonymous):

oops. a is correct

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

for #1, use wolfram alpha to get http://www.wolframalpha.com/input/?i=x^4-5*x^3%2B11*x^2-25*x%2B30+%3D+0 so from that page, the 4 answers are what you see attached